OpenStudy (anonymous):

A can of Coke is supposed to contain, on the average, 16 ounces of soda with a standard deviation of 0.4 ounces. Suspecting fraud, you take a random sample of 50 cans and measure the amount of Coke in each. Your measurements show that the 50 cans had a means 15.9 ounces. What is the probability of a random sample of 50 cans having a mean of 15.9 ounces of soda or less?

OpenStudy (anonymous):

is this the 1 you want help on?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright, lets see what we got...

OpenStudy (anonymous):

what grade is this?

OpenStudy (anonymous):

12th

OpenStudy (anonymous):

alright thanks. I believe we need to find the z score, and use a z table, do you have those

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

rad, ok, calculate the z score then with the numbers they gave us

OpenStudy (anonymous):

calculate 16+50+15.9=71.90

OpenStudy (anonymous):

i was thinking this equation \[z = \frac{ x-\upsilon }{ \sigma }\]

OpenStudy (anonymous):

x = the mean of sample data set u = the mean of ALL data sigma = the standard deveiation

OpenStudy (anonymous):

z=16-71.9/0.4

OpenStudy (anonymous):

or is it 16-15.9/0.4

OpenStudy (anonymous):

16-15.9/0.4 this one

OpenStudy (anonymous):

wait, the 16 and 15.9 are backwards its (15.9-16)/0.4

OpenStudy (anonymous):

-24.1

OpenStudy (anonymous):

close, in math you have to do ( ) then divide do this first: (15.9-16)

OpenStudy (anonymous):

(15.9-16)=-0.1

OpenStudy (anonymous):

yep, then do -0.1/0.4

OpenStudy (anonymous):

sorry (15.9-16)/0.4=0.25

OpenStudy (anonymous):

(15.9-16)/0.4=-0.25

OpenStudy (anonymous):

good, now look on your z table and find the value for -0.25

OpenStudy (anonymous):

I met -0.25

OpenStudy (anonymous):

find the probability by taking you z score value of -0.25 and taking the numbers like this on the graph